### v1

@@ Abstract
Affective Behavior Analysis aims to develop emotionally intelligent technology that can recognize and respond to human emotions. To advance this, the 7th Affective Behavior Analysis in-the-wild (ABAW) competition establishes two tracks: i.e., the Multi-task Learning (MTL) Challenge and the Compound Expression (CE) challenge based on Aff-Wild2 and C-EXPR-DB datasets. In this paper, we present our methods and experimental results for the two competition tracks. Specifically, it can be summarized in the following four aspects: 1) To attain high-quality facial features, we train a Masked-Auto Encoder in a self-supervised manner. 2) We devise a temporal convergence module to capture the temporal information between video frames and explore the impact of window size and sequence length on each sub-task. 3) To facilitate the joint optimization of various sub-tasks, we explore the impact of sub-task joint training and feature fusion from individual tasks on each task performance improvement. 4) We utilize curriculum learning to transition the model from recognizing single expressions to recognizing compound expressions, thereby improving the accuracy of compound expression recognition. Extensive experiments demonstrate the superiority of our designs.
@@ 1 Introduction
To facilitate this, the 7th Affective Behavior Analysis competition (ABAW7) set two competition tracks, i.e., the Multi-task Learning (MTL) Challenge and the Compound Expression (CE) [4, 10, 41, 45] Challenge based on the Aff-Wild [22] and C-EXPR-DB [13] datasets. Specifically, MTL actually involves three sub-tasks, i.e. the action unit (AU) prediction, the expression recognition (EXPR) [27, 38, 45, 5, 57], and the valence-arousal (VA) estimation [24, 12, 18, 30, 34]. In this track, participants are encouraged to leverage the commonalities and differences across these tasks. To improve the final performance of the model, we focus on three aspects: i) Enabling the model to learn general and powerful features that can be applied across tasks, enhancing the overall capability. ii) Achieving knowledge transfer between tasks, so features effective for one task can benefit others. iii) Fully utilizing the spatiotemporal features in the data to further boost model performance.
We took the following tries to achieve our goal. We integrate almost all the existing facial datasets and train a feature extractor in a self-supervised manner. This extractor can be applied to all tasks to obtain strong features. To fully utilize the facial expression features captured by different models, we first train models for each task individually. Then we devise a feature fusion module to integrate the features extracted by these pre-trained models with the features of the current training model. Moreover, we employ a Temporal Convergence module to capture and integrate dynamic changes in expressions, fully exploiting the temporal information in the dataset.
For the CE challenge, participations are required to achieve compound expression recognition in videos with limited amounts and unknown labels [4, 10, 41, 45]. We summarize the challenges into the following two aspects: i) How to utilize existing data to train models effectively and validate their performance on compound expressions with these labeled data. ii) Compared to basic single expressions, compound expressions are more complex and nuanced. Therefore, the second goal is to enable models to accurately recognize and capture the subtle changes in compound expressions.
To achieve these goals, we first utilize existing expression recognition datasets to train the model. Then we employ CutMix and Mixup techniques to generate compound expressions, facilitating the model to better understand complex emotions. Additionally, we introduce curriculum learning to transition from single expression recognition tasks to compound expression recognition. This gradual increase in task difficulty allows the model to adapt more stably to complexity, enhancing its generalization ability and leading to better performance on unseen data. It also efficiently utilizes data by first training on simpler datasets and then introducing more complex ones, maximizing the value of all available data.
We conduct comprehensive evaluations on the officially provided dataset and our self-curated validation dataset. The results demonstrate the effectiveness of our method design.
@@ 2 Related Work
2.1 Multi-task Learning for Affective Behaviour Analysis
@@
VA estimation involves predicting continuous emotional dimensions and has been traditionally approached using single-task models. Recent MTL approaches, such as Kollias et al. [21], have demonstrated that learning VA jointly with other tasks like ER and AU detection can enhance performance by capturing shared emotional cues. Their work utilized deep neural networks to jointly learn VA, ER, and AU, leading to improved generalizability. Expression recognition has significantly benefited from MTL frameworks. Zhang et al. [56, 50] introduce an MTL approach that jointly learns ER with VA estimation and AU detection. Their model utilizes a shared feature extractor with task-specific heads, resulting in better accuracy and robustness across tasks. They highlighted the importance of capturing the interdependencies between expressions, valence, and arousal. AU detection is another area where MTL has been effectively applied. Zhang et al. [53, 54] propose a framework that simultaneously performs AU detection and ER, leveraging the correlations between AUs and expressions.
@@
2.2 Compound Expression Recognition
Compound Expression Recognition (CER) extends traditional emotion recognition by identifying complex, blended expressions. Early efforts by [4, 10] systematically categorized compound emotions, laying the groundwork for further studies. With the rise of deep learning, methods utilizing convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have shown significant improvements in accuracy [8, 11, 3, 48, 36, 35]. For instance, Hasani and Mahoor [9] employ deep residual networks for CER, while Kosti et al. [26] use LSTM networks to capture temporal dynamics in video sequences. Multi-task learning (MTL) frameworks, as explored by Zhang et al. [51], have also enhanced CER by leveraging shared feature extractors with task-specific heads. Datasets like the Compound Facial Expressions of Emotion (CFEE) and RAF-DB have been instrumental in advancing CER research. Despite progress, challenges remain due to the subtlety and variability of compound expressions, highlighting the need for more sophisticated models and diverse datasets.
@@ 3 Method
In this work, we employ Masked-Auto Encoder as our feature extractor as it In this section, we present our approach for the two competition tracks in three parts. First, we will describe the construction process of the feature extractor utilized for both tracks. Then we introduce our method design for the Multi-Task Learning (MTL) challenge. In the final part, we detail our framework flow for the Compound Expression Recognition (CER) challenge.
Figure 1: Illustration of our proposed frameworks for the MTL and CE challenges. In the MTL track, we fuse features from other task models with the current model’s features and utilize the Temporal Convergence Module to capture dynamic expression information. The current task can also perform joint training with other tasks by integrating their classification heads. In the CE track, we start by training a multi-class model with single expression labels. In later stages, we add compound expression data, gradually increasing its proportion. These compound expressions come from both real data and data augmentation.
3.1 Feature Extractor
In this competition, we employ Masked-Auto Encoder as our feature extractor due to its powerful feature extraction capabilities. To obtain high-quality facial features, we first integrate a large-scale facial dataset, including AffectNet [32], CASIA-WebFace [46], CelebA [31], IMDB-WIKI [40], and WebFace260M [59]. The data undergoes further cleaning by the annotation platform to ensure its quality, and ultimately, 4.5 million images are used in the training process. Based on this dataset, we train the MAE in a self-supervised manner. Specifically, we randomly mask 75% of each facial image and then reconstruct the original face image. During the training stage, we utilize L2 Loss to calculate the difference between the reconstructed image and the original image for model optimization. To better adapt the MAE model to the affective behavior analysis, we fine-tuned the model on the AffectNet and data from Aff-Wild provided by the competition.
@@ 3.2 Multi-Task Learning
In the MTL track, as depicted in Fig. 1, we adopt a progressive learning approach, transitioning from single-task learning to multi-task joint learning. Initially, we construct classification or regression heads for the VA, ER, and AU tasks behind the MAE feature extractor and train them on the respective task’s effective data (i.e., excluding data with VA labels of -5, and expression and AU labels of -1). Next, we explore the impact of joint training by combining various tasks. We employ the feature extractor with multiple heads for the joint training of different tasks. Additionally, we investigate the effect of integrating features from different sub-tasks into other tasks on model training performance. Specifically, we first extract the output features of the MAE for each task after individual training. When training other tasks, we introduce a feature fusion module behind the MAE to combine the features of other tasks with those of the task being trained. We provide further results in the experimental section demonstrating the performance improvements from the progressive learning strategy.
@@
3.3 Compound Expression Recognition
3.3.1 Method Overview.
In our approach, we begin by training the model on single expression recognition tasks. This involves learning to identify basic emotions such as Neutral, Anger, Disgust, Fear, Happiness, Sadness, and Surprise. Once the model achieves satisfactory performance on the single expression recognition task, we introduce compound expression recognition. This progressive training strategy increases the difficulty of tasks. It allows the model to adapt to the more complex task of identifying blended or compound expressions. Moreover, in the second learning stage, we utilize Cutmix and Mixup data augmentation techniques to generate compound expressions. This increases dataset diversity, enhancing the robustness and generalization of our model.
3.3.2 Single Expression Recognition.
To attain the high-quality feature extractor, we first finetune the MAE model on RAF-DB and our curated Fuxi-EXPR dataset (including 25K images with seven basic expressions as RAF-DB [28]). Note that, in this stage, we just utilize the data with the single expression label to train the model, as shown in Fig. 1. To increase the diversity of the training dataset, we adopt the cutout data augmentation technique as well as the basic data augmentation techniques such as random flipping, color jittering, and random crop in the initial learning stage. Note that considering the final compound expressions do not include the neutral label, we filter out all data with the neutral label from the entire dataset.
3.3.3 Compound Expression Recognition.
As the model’s performance stabilizes, we progressively introduce more complex compound expressions to train our model. To reduce the risk of overwhelming the model with difficult data too early, we progressively increase the proportion of compound expressions in the training data. Specifically, we divide the overall training stage into four sub-stages, with the number of training epochs for each stage being 5, 5, 3, and 3. The proportion of compound expressions in each stage is 0, 0.2, 0.4, and 1, respectively. During the training phase, we treat the compound expression recognition task as a multi-class classification problem. To achieve better performance, during the inference phase, we constrain the final results into the officially provided compound expression categories. Specifically, we sum the classification probabilities of the individual expression categories involved in the compound expressions and utilize the argmax of these summed probabilities as the final result.
@@ 3.4 Training Objectives
3.4.2 Objectives for Compound Expression Recognition.
In the whole training stage, we utilize Binary Cross-Entropy Multi-Label Loss as the optimization objective, mathematically expressed by:
$\mathcal{L}_{BMC}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}\left[y_{ic}\log(p_{ic})+(1-y_{ic})\log(1-p_{ic})\right].$ (7)
Here, $N$ is the number of samples, $C$ is the number of categories, $y_{ic}$ is the binary indicator (0 or 1) if class label $c$ is the correct classification for sample $i$, and $p_{ic}$ is the predicted probability of sample $i$ being of class $c$.
@@ 4 Experiments
4.1.2 Data for CE.
For the CE track, the organizers provide 56 video data from the C-EXPR-DB [13] database. These videos are categorized into the following categories: Fearfully Surprised, Happily Surprised, Sadly Surprised, Disgustedly Surprised, Angrily Surprised, Sadly Fearful, and Sadly Angry. Note that we do not have the ground truth labels for these videos.
@@
Except for the official provided data, we also utilize the images from AffectNet [32], CASIA-WebFace [46], CelebA [31], IMDB-WIKI [40], WebFace260M [59], and the Fuxi-EXPR private dataset. Fuxi-EXPR contains 71,618 facial images collected from films, television works, and public video platforms. Data cleaning and management are handled by the Netease Fuxi Youling Crowdsourcing platform [^2] https://fuxi.163.com/solution/data. The dataset includes labels for six basic emotions (Surprised, Fear, Disgust, Happiness, Sadness, Anger) and an additional category for Others. For the AU detection track, we supplement the Aff-wild2 dataset with the BP4D dataset [55] to address the limited coverage of certain AU categories. For the expression recognition track, we enhance the dataset with RAF-DB [28] and AffectNet [32] to increase the data of Anger, Disgust, and Fear expressions. For the CE track, we utilize RAF-DB and Fuxi-EXPR as the training and validation datasets. To obtain stable performance, we retain only the six basic emotions for the two datasets.
@@ 4.2 Metrics
4.2.2 Metric for CE.
For the CE track, the performance $P$ is measured by the average F1 score across all seven categories, expressed by:
$P=\frac{\sum_{c=1}^{7}F1_{c}}{7}.$ (9)
4.3 Implementation Details.
All training images are resized to 224 $\times$ 224 pixels. The MAE pertaining is conducted on 8 NVIDIA A30 GPUs with a batch size of 4096. For training details of MTL and CE tracks, all training configurations (including the batch size, optimizer, scheduler, and learning rate et al.) can be found in our released code. All experiments are conducted on 4 NVIDIA A100 GPUs.
4.4 Experimental Results.
@@
We evaluate our final approach to the MTL challenge on the official validation set. Since the CE challenge does not involve the validation set, we evaluate our method to the CE challenge on the validation set of RAF-DB-b and FUXI-EXPR. The results are shown in Table 1.
Table 1: Quantitative results on the officially provided validation dataset. Note that we obtain the results of MTL by combining the results from various models.
@@
CE _ _ _ _ _ 0.6856
In the MTL track, our method significantly outperformed the baseline model, achieving scores of 0.6628, 0.4883, and 0.6190 in Average CCC, $F_{expr}$, and $F_{aus}$, respectively. Note that our results for the MTL task are from combining the best models for each sub-task. For the CE task, we achieved an F-score of 0.6856.
@@
Table 2: Impact of the Feature Fusion Module. All experiments are conducted based on independently trained models. Here, $Per.$ represents the performance of each model on its respective task.
@@
As suggested in Table 2, combining features from other task models via the feature fusion module does not universally enhance the performance of all task models. The degree of performance improvement varies significantly depending on the fused features. Specifically, for the AU task model, the best performance (F-score of 0.5001) is achieved without using the feature fusion module, and introducing features from other task models decreases performance by 0.28% to 1.75%. Conversely, for the EXPR task model, fusing features from the AU task model improves performance by 1.4%. For the Valence task model, the best results come from incorporating features from both the AU and Arousal models, resulting in a 1.74% improvement. For the Arousal task, integrating features from both the AU and Valence models leads to a significant performance increase of 5.69%.
@@
Table 4: Analysis of Data Augmentation. All experiments are conducted under the optimal setting of Epoch Distribution and Compound Data Proportion Distribution. The blended weight of the Mixup is set to 0.1.
Exp Mixup CutMix P
1 0.6764
2 0.6722
3 0.6856
@@
For the EXPR task model, the highest F-score of 0.4883 is achieved by fusing features from the AU and Valence models and jointly training with the AU, EXPR, and Valence estimation tasks. For the Valence estimation task, the best performance, with a CCC value of 0.6628, is obtained by combining features from the AU and Valence models and training under the Valence, Arousal estimation, and AU prediction tasks. For the Arousal estimation task, the optimal result is achieved by incorporating features from the AU and Valence models and jointly training with the AU, Valence, and Arousal estimation tasks.
@@
30 30 0.6061
Table 6: Analysis of Curriculum Learning. Epoch Dis. (Epoch Distribution) without numerical values indicates training with only single-label data. Here, Epoch Dis. indicates the epoch number of each training stage. Compound Prop (Compound Data Proportion Distribution) is the proportion of the compound data in each training stage.
       Exp        Epoch Dis.        Compound Prop.        P
       1        _        _        0.6653
       2        [5, 15]        [0,1]        0.6759
       3        [5, 5, 5]        [0,0.5,1]        0.6652
       4        [5, 5, 3, 3]        [0,0.2,0.4,1]        0.6856
       5        [5, 3, 3, 3, 3]        [0,0.2,0.4,0.6,1]        0.6325
4.4.5 Analysis of Curriculum Learning.
We evaluate our CE recognition model with different epoch distributions and compound data proportion distributions. The results are presented in Table 6. As indicated in Table 6, the best performance is achieved under the Epoch distribution of [5, 5, 3, 3] and Compound Propotion distribution of [0, 0.2, 0.4, 1], with the F1 score of 0.6856. Moreover, we found that the model performance does not always improve with an increasing number of training stages. When we divide the training process into five stages and adjust the distribution of compound data to be more even across these stages, the performance decreased by 5.31% compared to the four-stage training process.
4.4.6 Analysis of Data Augmentation.
We evaluate the model performance when incorporating different data augmentation methods in the training stage. As shown in Table 4, the introduction of Mixup and CutMix methods both enhance the model performance. Incorporating both data augmentation methods simultaneously achieves the highest F1-score value.

### v3

@@ Abstract
Affective Behavior Analysis aims to develop emotionally intelligent technology that can recognize and respond to human emotions. To advance this field, the 7th Affective Behavior Analysis in-the-wild (ABAW) competition holds the Multi-Task Learning Challenge based on the s-Aff-Wild2 database. The participants are required to develop a framework that achieves Valence-Arousal Estimation, Expression Recognition, and AU detection simultaneously. To achieve this goal, we propose a progressive multi-task learning framework that fully leverages the distinct focuses of each task on facial emotion features. Specifically, our method design can be summarized into three main aspects: 1) Separate Training and Joint Training: We first train each task model separately and then perform joint training based on the pre-trained models, fully utilizing the feature focus aspects of each task to improve the overall framework performance. 2) Feature Fusion and Temporal Modeling: We investigate effective strategies for fusing features extracted from each task-specific model and incorporate temporal feature modeling during the joint training phase, which further refines the performance of each task. 3) Joint Training Strategy Optimization: To identify the optimal joint training approach, we conduct a comprehensive strategy search, experimenting with various task combinations and training methodologies to further elevate the overall performance of each task. According to the official results, our team achieves first place in the MTL challenge with a total score of 1.5286 (i.e., AU F-score 0.5580, Expression F-score 0.4286, CCC VA score 0.5420). Our code is publicly available here.
@@ 1 Introduction
To facilitate this, the 7th Affective Behavior Analysis competition (ABAW7) [24] set two competition tracks, i.e., the Multi-task Learning (MTL) Challenge and the Compound Expression (CE) [4, 9, 43, 48] Challenge based on the Aff-Wild [20] and C-EXPR-DB [11] datasets. Specifically, MTL actually involves three sub-tasks, i.e. the action unit (AU) prediction, the expression recognition (EXPR) [25, 39, 48, 5, 58], and the valence-arousal (VA) estimation [22, 10, 16, 31, 35]. In this track, participants are encouraged to leverage the commonalities and differences across these tasks. To improve the final performance of the model, we focus on three aspects: i) Enabling the model to learn general and powerful features that can be applied across tasks [28, 29, 55, 36, 37], enhancing the overall capability. ii) Achieving knowledge transfer between tasks, so features effective for one task can benefit others. iii) Fully utilizing the spatiotemporal features in the data to further boost model performance.
To be specific, we first integrate almost all the existing facial datasets and train a feature extractor in a self-supervised manner. This extractor can be applied to all tasks to obtain strong and representative features. To fully utilize the facial expression features captured by different models, we first train models for each task individually. Then we devise a feature fusion module to integrate the features extracted by these pre-trained models with the features of the current training model. Moreover, we employ a Temporal Convergence module to capture and integrate dynamic changes in expressions, fully exploiting the temporal information in the dataset. Comprehensive evaluations conducted on the officially provided validation and test datasets the superiority of our method.
@@ 2 Related Work
VA estimation involves predicting continuous emotional dimensions and has been traditionally approached using single-task models. Recent MTL approaches [19] have demonstrated that learning VA jointly with other tasks like ER and AU detection can enhance performance by capturing shared emotional cues. Their work utilized deep neural networks to jointly learn VA, ER, and AU, leading to improved generalizability. Expression recognition has significantly benefited from MTL frameworks. Zhang et al. [57, 52] introduce an MTL approach that jointly learns ER with VA estimation and AU detection. Their model utilizes a shared feature extractor with task-specific heads, resulting in better accuracy and robustness across tasks. They highlighted the importance of capturing the interdependencies between expressions, valence, and arousal. AU detection is another area where MTL has been effectively applied. Zhang et al. [54, 55] propose a framework that simultaneously performs AU detection and ER, leveraging the correlations between AUs and expressions.
@@ 3 Method
Figure 1: Illustration of our proposed framework for the MTL competition. In the initial stage, we train models for each task separately. Once optimal performance is achieved, we begin joint training to enhance performance further. Taking EXPR model training as an example, features like $F_{AU}$, $F_{V}$, and $F_{A}$ from pre-trained encoders are fused with features from the current EXPR Encoder via the Feature Fusion Model. The fused features are then processed by the Temporal Convergence Module to capture temporal information. Finally, the features are sent to task-specific heads (AU, EXPR, and VA) for predictions. The final integration scheme for each subtask depends on the validation set performance, and the detailed analysis is provided in Sec. 4
@@ 3.1 Multi-Task Learning
In the MTL track, as depicted in Fig. 1, we adopt a progressive learning approach, transitioning from single-task learning to multi-task joint learning. Initially, we construct classification or regression heads for the VA, ER, and AU tasks behind the MAE feature extractor and train them on the respective task’s effective data (i.e., excluding data with VA labels of -5, and expression and AU labels of -1). Next, we explore the impact of joint training by combining various tasks. We employ the MAE as our feature extractor [8, 55, 30] for all sub-tasks. Additionally, we investigate the effect of integrating features from different sub-tasks into other tasks on model training performance. Specifically, we first extract the output features of the MAE for each task after individual training. When training other tasks, we introduce a feature fusion module behind the MAE to combine the features of other tasks with those of the task being trained. We provide further results in the experimental section demonstrating the performance improvements from the progressive learning strategy.
@@ 4.1 Dataset
In addition to the official data, we utilize the images from AffectNet [33], CASIA-WebFace [49], CelebA [32], IMDB-WIKI [41], WebFace260M [60]. We train the initial feature extractor on these datasets using a facial reconstruction task. After that, we fine-tune the extractor on BPD4 [56] by predicting action units and then use this trained model as the initial feature extractor for all tasks.
@@ 4.2 Metrics
4.3 Implementation Details
All training images are resized to 224 $\times$ 224 pixels. The MAE pertaining is conducted on 8 NVIDIA A30 GPUs with a batch size of 4096. For training details of the multi-task learning track, all training configurations (including the batch size, optimizer, scheduler, and learning rate et al.) can be found in our released code. All experiments are conducted on 4 NVIDIA A100 GPUs.
4.4 Experimental Results
@@
We evaluate our final approach to the MTL challenge on the official validation set. Since the CE challenge does not involve the validation set, we evaluate our method to the CE challenge on the validation set of RAF-DB-b and FUXI-EXPR. Results are shown in Table 1.
Table 1: Quantitative results on the officially provided validation dataset.
@@
Table 2: Impact of the Feature Fusion Module. All experiments are conducted based on separately trained models. Here, $Per.$ represents the performance of each model on its respective task.
@@
As suggested in Table 2, combining features from other task models via the feature fusion module does not universally enhance the performance of all task models. The degree of performance improvement varies significantly depending on the fused features. Specifically, for the AU task model, the best performance (F-score of 0.5001) is achieved without using the feature fusion module, and introducing features from other task models decreases performance by 0.28% to 1.75%. Conversely, for the EXPR task model, fusing features from the AU task model improves performance by 1.4%. For the Valence task model, the best results come from incorporating features from both AU and Arousal models, resulting in a 1.74% improvement. For the Arousal task, integrating features from both the AU and Valence models leads to a significant performance increase of 5.69%.
@@
For the EXPR task model, the highest F-score of 0.5030 is achieved by fusing features from the AU and Valence models and jointly training with the AU, EXPR, and Valence estimation tasks. For the Valence estimation task, the best performance, with a CCC value of 0.6926, is obtained by combining features from the AU and Valence models and training under the Valence, Arousal estimation, and AU prediction tasks. For the Arousal estimation task, the optimal result is achieved by incorporating features from the AU and Valence models and jointly training with the AU, Valence, and Arousal estimation tasks.
@@
30 30 0.6139
4.5 Evaluation on the Test Dataset
In Table 5, we show the final evaluations on the official test dataset. Our team (i.e. Netease Fuxi AI Lab) attains a first place in the Multi-Task Learning competition. Overall, our framework outperforms the second-place team (i.e., HSEomotion) by approximately 28.14%. Specifically, our method demonstrates exceptional performance in the VA Estimation task, surpassing the second-place team by nearly 14%. Additionally, in the AU Detection and Expression Recognition tasks, our method outperforms the second-place team by 4.61% and 9.89%, respectively. This fully proves the effectiveness of our solution. As suggested by the results of all teams, the F1 score for the Expression Recognition task was relatively lower compared to the other two tracks. We attribute this to the limited amount of training data. The amount of the training data for Expression Recognition is one-third less than that for the other two tracks, while the amount of validation and test data is roughly the same across all three tracks.
Table 5: There are the final results on the official test dataset. Our team ( i.e. Netease Fuxi AI Lab) attains the best results on MTL, outperforming the second place by 28.14% in the overall metric. Moreover, our method also shows outstanding performance in the AU, EXPR, and VA tasks.
Teams Overall Metric AUs F1 Score EXPR F1 Score CCC VA
baseline [24] 0.3400 0.1200 0.1007 0.1193
SML 0.8692 0.4046 0.1938 0.4074
AIWELL-UOC [3] 1.1145 0.4663 0.2772 0.3783
SCU ACers [26] 1.1640 0.4879 0.3018 0.3743
HFUT-MAC1 [44] 1.1777 0.4997 0.2997 0.3783
HSEmotion [42] 1.2472 0.5119 0.3279 0.4074
Netease Fuxi AI Lab [30] 1.5286 0.5580 0.4286 0.5420